1.      The velocity of a train changes from 0 to 60 in 20 seconds. Assuming that it is uniformly accelerated, what is the acceleration in m/s squared? 


1
Expert's answer
2021-11-10T17:50:03-0500

u=0m/sec

v=60m/sec

t=20m/sec

v=u+at

"a=\frac{v-u}{t}"

Put value

"a=\frac{60-0}{20}=3m/sec^2"